# Vulnerability Management Chaining: An Integrated Framework for Efficient Cybersecurity Risk Prioritization

### v2

@@ Abstract
Cybersecurity teams face an overwhelming vulnerability crisis: with 25,000+ new CVEs disclosed annually, traditional CVSS-based prioritization requires addressing approximately 57% of all vulnerabilities while correctly identifying only 20% of those actually exploited. We propose Vulnerability Management Chaining, an integrated decision tree framework combining historical exploitation evidence (KEV), predictive threat modeling (EPSS), and technical impact assessment (CVSS) to transform vulnerability management from reactive patching to strategic threat-driven prioritization. Experimental validation using 28,377 real-world vulnerabilities demonstrates approximately 14-18 fold efficiency improvements while maintaining 85+ percent coverage of actual threats. Organizations can reduce urgent remediation workload by approximately 95% (from  16,000 to  850 vulnerabilities). The integration identifies 57 additional exploited vulnerabilities that neither KEV nor EPSS captures individually. Our framework uses exclusively open-source data, democratizing advanced vulnerability management regardless of budget or expertise. This research establishes the first empirically validated methodology for systematic vulnerability management integration, with immediate applicability across diverse organizational contexts.
@@ 1.2 Limitations of Current Approaches
The Common Vulnerability Scoring System (CVSS) has served as the de facto standard for vulnerability management since its introduction. CVSS provides a standardized method for rating the severity of security vulnerabilities based on their technical characteristics, including attack complexity, required privileges, and potential impact on confidentiality, integrity, and availability.
However, CVSS-based vulnerability management has come under increasing criticism for its fundamental inefficiency. Research has consistently shown that only a small fraction of high-CVSS vulnerabilities are ever exploited in practice. Studies indicate that fewer than 20% of vulnerabilities rated as "Critical" or "High" by CVSS are actually observed being exploited in real-world attacks [2], yet these high-severity ratings are assigned to approximately 57% of all published vulnerabilities [3].
This mismatch creates a significant operational burden, leading researchers to question whether it’s "time to change the CVSS" [3] and describing the system as "ubiquitous and broken" [4]. Security teams following CVSS-based prioritization must address thousands of vulnerabilities that pose little real-world risk. Simultaneously, they may overlook vulnerabilities that attackers actively exploit but receive lower CVSS scores due to complex attack vectors or authentication requirements.
Government agencies have recognized these limitations, with CISA implementing binding operational directives that mandate specific remediation timelines [5, 6]. These directives represent a shift toward evidence-based vulnerability management, requiring federal agencies to prioritize vulnerabilities with confirmed exploitation evidence regardless of their CVSS scores.
Furthermore, CVSS was designed to measure technical severity rather than risk in the broader sense. The Forum of Incident Response and Security Teams (FIRST), which maintains CVSS, explicitly states that CVSS base scores should not be used alone for risk assessment and recommends incorporating additional contextual factors [7]. However, in practice, most organizations rely primarily on CVSS base scores due to the complexity and resource requirements of implementing more comprehensive risk assessment frameworks.
1.3 Emerging Threat Intelligence Approaches
Recognizing the limitations of traditional vulnerability management, several threat intelligence-based approaches have emerged to provide more effective prioritization. In 2021, the U.S. Cybersecurity and Infrastructure Security Agency (CISA) introduced the KEV catalog, which identifies vulnerabilities that have been observed being exploited in real-world attacks. KEV represents a paradigm shift from theoretical severity assessment to evidence-based prioritization, requiring federal agencies to remediate KEV-listed vulnerabilities within specified timeframes regardless of their CVSS scores.
Developed by researchers and maintained by FIRST, EPSS uses machine learning to predict the probability that a vulnerability will be exploited within the next 30 days. EPSS incorporates numerous data sources, including threat intelligence feeds, proof-of-concept availability, and social media discussions, to generate probabilistic exploitation scores.
While both KEV and EPSS represent significant advances over pure CVSS-based prioritization, each has important limitations. KEV provides high-confidence information about past exploitation but cannot predict future threats. EPSS offers forward-looking predictions but may miss vulnerabilities that are already being exploited. Additionally, EPSS is explicitly not recommended for use as a standalone risk assessment tool, particularly for vulnerabilities where exploitation has already been confirmed.
@@ 1.5 Proposed Approach and Contributions
Our primary contributions include an integrated framework design that develops a practical decision tree framework systematically combining three major vulnerability management approaches, addressing the limitations of using any single method in isolation. We provide empirical validation by demonstrating the effectiveness of our approach using real-world data including 28,377 CVEs published over a 13-month period, actual network intrusion detection data, and public security vendor reports of exploited vulnerabilities.
@@ 1.6 Paper Organization
The remainder of this paper is structured as follows. Section 2 provides detailed background on existing vulnerability management approaches and establishes the theoretical foundation for our integrated methodology. Section 3 presents our Vulnerability Management Chaining framework, including the decision tree design and implementation considerations. Section 4 describes our experimental methodology and presents comprehensive evaluation results. Section 6 discusses the implications of our findings, limitations of the current approach, and directions for future work. Section 7 concludes with a summary of contributions and their broader significance for the cybersecurity community.
@@ 2 Background and Related Work
This section provides essential background on the three vulnerability management frameworks that form the foundation of our approach, reviews related work in vulnerability prioritization, and identifies the specific gaps that motivate our research.
@@ 2.1 Common Vulnerability Scoring System (CVSS)
CVSS, maintained by the Forum of Incident Response and Security Teams (FIRST), provides a standardized framework for rating the severity of security vulnerabilities [8]. The system generates numerical scores from 0.0 to 10.0 based on eight metrics that capture the fundamental characteristics of vulnerabilities.
CVSS Base Score calculations consider two primary factors: exploitability metrics including Attack Vector, Attack Complexity, Privileges Required, and User Interaction, along with impact metrics covering Confidentiality, Integrity, Availability impacts, plus Scope. For example, a remotely exploitable vulnerability requiring no authentication that completely compromises system confidentiality, integrity, and availability would receive the maximum score of 10.0.
The scoring system categorizes vulnerabilities into severity levels: Critical (9.0-10.0), High (7.0-8.9), Medium (4.0-6.9), and Low (0.1-3.9). Government agencies and industry standards frequently mandate remediation timeframes based on these categories, with critical vulnerabilities often requiring patching within 15 days and high-severity vulnerabilities within 30 days [5].
However, research has consistently demonstrated a significant disconnect between CVSS scores and real-world exploitation patterns. Analysis of vulnerability databases shows that approximately 57% of published vulnerabilities receive High or Critical CVSS ratings, yet studies of actual exploitation patterns indicate that fewer than 20% of these highly-rated vulnerabilities are ever observed being exploited [9, 10].
@@ 2.2 Exploit Prediction Scoring System (EPSS)
EPSS addresses CVSS limitations by using machine learning to predict the probability that a vulnerability will be exploited within 30 days [11]. The system is maintained by FIRST [12] and represents a significant advancement in predictive vulnerability assessment. [^1] This study utilizes EPSS v3.0. During manuscript preparation, EPSS v4.0 was released with enhanced predictive capabilities, and alternative predictive metrics such as Likely Exploited Vulnerabilities (LEV) have emerged [13], further validating the evolution toward sophisticated threat prediction methodologies. The current EPSS model (version 3.0) incorporates over 1,400 features including vulnerability characteristics, proof-of-concept availability, references in security advisories, social media mentions, and exploitation data from security vendors.
EPSS scores range from 0 to 1, representing the probability of exploitation. The system is updated daily, allowing scores to evolve as new information becomes available. Research shows that EPSS significantly outperforms CVSS for identifying vulnerabilities likely to be exploited, achieving 82% coverage of exploited vulnerabilities while requiring remediation of only 45.5% of the total vulnerability population, compared to CVSS which achieves similar coverage but requires addressing 96.1% of vulnerabilities [11].
However, EPSS has important limitations. The system is explicitly not recommended as a standalone risk assessment tool, as it focuses solely on threat likelihood without considering potential impact or organizational context. Additionally, EPSS is not designed for vulnerabilities where exploitation has already been confirmed, as the prediction model assumes unknown exploitation status. Recent enhancements to EPSS incorporate community-driven insights and expanded data sources [14], demonstrating continued evolution in predictive vulnerability assessment methodologies.
2.3 Known Exploited Vulnerabilities (KEV)
In response to high-profile attacks exploiting known vulnerabilities, CISA introduced the KEV catalog in 2021 [15]. KEV lists vulnerabilities that meet three criteria: assigned CVE identifier, confirmed evidence of active exploitation, and available vendor-provided remediation. As of 2023, the catalog contains over 900 vulnerabilities spanning multiple years.
KEV represents a significant policy shift toward evidence-based vulnerability management. Binding Operational Directive 22-01 requires federal agencies to remediate KEV-listed vulnerabilities within specified timeframes (typically 14 days) regardless of CVSS scores [6]. This approach acknowledges that confirmed exploitation provides stronger evidence of risk than theoretical severity assessments.
While KEV provides high-confidence threat intelligence, it has notable limitations. The catalog only includes vulnerabilities with confirmed exploitation evidence, potentially missing emerging threats that have not yet been widely observed. Additionally, KEV focuses primarily on vulnerabilities affecting federal networks, which may not fully represent threat landscapes for other sectors.
2.4 Related Work in Vulnerability Management
Several research efforts have attempted to improve vulnerability prioritization beyond traditional CVSS-based approaches. Commercial security vendors have developed proprietary risk-scoring systems that combine vulnerability severity with threat intelligence and environmental context [16, 17]. Alternative academic approaches have explored offensive security perspectives for vulnerability prioritization [18]. While these systems show promise, they typically require expensive subscriptions and use proprietary algorithms that limit transparency and reproducibility.
Various researchers have applied machine learning to vulnerability prioritization [19, 20]. These approaches often achieve good performance on specific datasets but face challenges with generalizability and the need for extensive labeled training data. Some research has applied formal decision analysis techniques to vulnerability prioritization [21, 22]. While mathematically rigorous, these approaches often require significant expertise to implement and may not scale to operational environments.
Developed by researchers at Carnegie Mellon’s CERT Coordination Center, SSVC provides decision trees for different stakeholder roles including suppliers, deployers, and coordinators [10]. The framework has been refined through multiple iterations [23], offering more nuanced decision-making than single numerical scores but requiring manual assessment of multiple factors for each vulnerability.
Various research efforts have explored integrating threat intelligence into vulnerability management. Studies have examined dark web monitoring for threat prediction [24] and proactive identification of emerging threats [25]. However, these approaches often require specialized data sources and expertise not available to all organizations.
2.5 Research Gaps and Motivation
Our analysis of existing approaches reveals several important gaps that motivate our research. While individual systems like CVSS, EPSS, and KEV each provide valuable information, no practical framework exists for systematically combining their insights. Organizations must manually reconcile conflicting signals from different systems. Advanced approaches like SSVC provide sophisticated decision-making frameworks but require significant expertise and time investment that may not be practical for many organizations.
Much existing research focuses on algorithmic improvements using synthetic or limited datasets, with insufficient validation using real-world operational data across diverse organizational contexts. Many promising approaches rely on proprietary data sources or commercial tools, limiting adoption by organizations with constrained security budgets.
Our research addresses these gaps by developing a practical integration framework that combines the strengths of existing open-source vulnerability management systems while providing empirical validation using real-world exploitation data. The resulting approach aims to be both theoretically sound and operationally practical for organizations of varying sizes and security maturity levels.
3 Proposed Methodology
@@
This section presents Vulnerability Management Chaining, our integrated framework that combines CVSS, EPSS, and KEV to achieve more efficient vulnerability prioritization while maintaining comprehensive coverage. We describe the theoretical foundation, decision tree design, and practical implementation considerations.
@@ 3.1 Framework Design Principles
Our approach aligns with established risk assessment frameworks [26] while addressing the practical limitations identified in vulnerability management research [27]. The design principles address specific shortcomings of existing approaches through a threat-first prioritization strategy that evaluates whether a vulnerability poses an active or likely threat before assessing technical characteristics. This approach focuses limited security resources on vulnerabilities with confirmed or predicted exploitation potential, rather than starting with technical severity assessment that may identify thousands of theoretically severe but practically irrelevant vulnerabilities.
The framework employs contextual refinement by applying additional criteria to assess attack feasibility and potential impact after identifying threat-relevant vulnerabilities. This enables more nuanced prioritization decisions that consider both the likelihood of exploitation and the practical constraints that may affect attack success or impact severity. Unlike complex alternatives that require proprietary data sources, our framework relies exclusively on freely available data sources including CVSS scores from NVD, EPSS scores from FIRST, and KEV listings from CISA, ensuring broad accessibility and reproducibility across organizations with varying security budgets.
The decision tree structure provides clear, actionable guidance without requiring complex mathematical calculations or subjective assessments that might introduce inconsistency across different analysts or organizational contexts. This operational simplicity enables high levels of automation while maintaining the sophistication necessary for effective threat intelligence integration.
3.2 Vulnerability Management Chaining Architecture
Figure 1 illustrates our complete decision tree framework, which processes vulnerabilities through two sequential evaluation stages that systematically combine threat likelihood assessment with vulnerability characteristic evaluation.
VulnerabilityAssessmentKEV ListedOREPSS $\geq$ 0.088CVSS $\geq$ 7.0High PriorityCVSS <<< 7.0Medium PriorityNot KEV ListedANDEPSS <<< 0.088Normal Process
Figure 1: Vulnerability Management Chaining decision tree framework
The first stage determines whether a vulnerability poses an active or imminent threat using two complementary data sources. Vulnerabilities listed in CISA’s KEV catalog receive immediate threat classification, as they have confirmed evidence of real-world exploitation. KEV provides high-confidence threat intelligence but only covers vulnerabilities where exploitation has been definitively observed and reported. For vulnerabilities not in KEV, we apply EPSS scores to identify those with high exploitation probability. Following established research [11, 14], we use a threshold of 0.088, which provides coverage comparable to CVSS 7.0+ while significantly reducing the number of vulnerabilities requiring priority attention.
The combination of KEV and EPSS addresses the limitations of either approach alone. KEV covers confirmed threats but cannot predict emerging exploitation, while EPSS provides forward-looking threat assessment but may miss currently exploited vulnerabilities. Together, they provide comprehensive threat coverage spanning both current and predicted exploitation scenarios.
Vulnerabilities identified as threat-relevant in the first stage undergo additional evaluation using CVSS base scores to assess attack feasibility and potential impact. High-impact vulnerabilities with CVSS scores of 7.0 or higher combine threat relevance with significant attack potential, warranting immediate attention. The CVSS 7.0 threshold aligns with widely adopted government and industry standards for "High" severity classification. Lower-impact vulnerabilities with CVSS scores below 7.0, while having confirmed or predicted exploitation potential, may warrant different treatment due to factors that reduce urgency such as complex attack prerequisites, limited impact scope, or significant attack complexity.
This two-stage approach enables organizations to maintain aggressive response timelines for the most critical vulnerabilities while applying more measured approaches to threats with mitigating factors. The threat-first architecture places threat metrics before vulnerability metrics because first filtering for threat-relevant vulnerabilities allows organizations to focus detailed analysis on a much smaller subset of the total vulnerability population, improving analytical efficiency. Confirmed or predicted exploitation provides stronger evidence for prioritization decisions than theoretical severity assessments alone, and as vulnerability disclosure rates continue to increase, threat-first filtering becomes increasingly important for maintaining manageable workloads.

### v3

@@ Abstract
As the number of Common Vulnerabilities and Exposures (CVE) continues to grow exponentially, security teams face increasingly difficult decisions about prioritization. Current approaches using Common Vulnerability Scoring System (CVSS) scores produce overwhelming volumes of high-priority vulnerabilities, while Exploit Prediction Scoring System (EPSS) and Known Exploited Vulnerabilities (KEV) catalog offer valuable but incomplete perspectives on actual exploitation risk. We present Vulnerability Management Chaining, a decision tree framework that systematically integrates these three approaches to achieve efficient vulnerability prioritization. Our framework employs a two-stage evaluation process: first applying threat-based filtering using KEV membership or EPSS threshold $\geq$ 0.088), then applying vulnerability severity assessment using CVSS scores $\geq$ 7.0) to enable informed deprioritization. Experimental validation using 28,377 real-world vulnerabilities and vendor-reported exploitation data demonstrates 18-fold efficiency improvements while maintaining 85.6% coverage. Organizations can reduce urgent remediation workload by approximately 95%. The integration identifies 48 additional exploited vulnerabilities that neither KEV nor EPSS captures individually. Our framework uses exclusively open-source data, enabling immediate adoption regardless of organizational resources.
@@ 1.2 Limitations of Current Approaches
The industry has long relied on the Common Vulnerability Scoring System (CVSS) as the primary method for vulnerability prioritization. However, CVSS was designed to measure the theoretical severity of vulnerabilities in isolation, not their likelihood of exploitation in real-world attacks. This fundamental limitation has led to a situation where organizations using CVSS-based prioritization face an overwhelming number of "high" and "critical" vulnerabilities, many of which are never actually exploited by attackers.
Recent efforts have introduced more sophisticated approaches to vulnerability management. The Exploit Prediction Scoring System (EPSS), developed by FIRST, uses machine learning to predict the likelihood of vulnerability exploitation within 30 days. While EPSS represents a significant advance in predictive threat intelligence, it remains a probabilistic model with inherent uncertainty and may miss vulnerabilities that become exploited through novel attack patterns.
The Cybersecurity and Infrastructure Security Agency (CISA) maintains the Known Exploited Vulnerabilities (KEV) catalog, which documents vulnerabilities with confirmed evidence of active exploitation. While KEV provides high-confidence threat intelligence, it is inherently retrospective and may not capture emerging threats or vulnerabilities exploited in targeted attacks that have not yet been publicly disclosed.
Each of these approaches captures important but incomplete aspects of vulnerability risk. CVSS measures potential impact but not exploitation likelihood. EPSS predicts exploitation probability but may have false positives and negatives. KEV confirms actual exploitation but only after attacks have occurred. Organizations need a systematic way to combine these complementary perspectives to achieve more effective vulnerability management.
@@ 1.4 Proposed Approach and Contributions
Our primary contributions include an integrated framework design that develops a practical decision tree framework systematically combining three major vulnerability management approaches, addressing the limitations of using any single method in isolation. We provide empirical validation by demonstrating the effectiveness of our approach using real-world data including 28,377 CVEs published over a 13-month period, and public security vendor reports of exploited vulnerabilities.
@@ 1.5 Paper Organization
The remainder of this paper is structured as follows. Section 2 provides background on vulnerability management systems and reviews related work. Section 3 presents our Vulnerability Management Chaining framework design and implementation details. Section 4 describes our experimental methodology including data collection and evaluation metrics. Section 5 presents comprehensive evaluation results comparing our approach with existing methods. Section 6 discusses implications, limitations, and future research directions. Section 7 concludes with key findings and practical recommendations.
@@ 2 Background and Related Work
This section provides essential background on current vulnerability management approaches and reviews related research that informs our integrated framework design.
@@ 2.1 Common Vulnerability Scoring System (CVSS)
The Common Vulnerability Scoring System represents the industry standard for vulnerability severity assessment. Maintained by the Forum of Incident Response and Security Teams (FIRST), CVSS provides a standardized method for capturing the principal characteristics of vulnerabilities and producing numerical severity scores ranging from 0.0 to 10.0 [2].
CVSS v3.1 calculates scores using three metric groups. Base metrics represent intrinsic vulnerability characteristics that remain constant over time and across environments, including attack vector, attack complexity, privileges required, user interaction requirements, and impact on confidentiality, integrity, and availability. Temporal metrics capture characteristics that change over time, such as exploit code maturity and available patches. Environmental metrics allow organizations to customize scores based on their specific deployment context.
Despite its widespread adoption, CVSS faces significant criticism regarding its effectiveness for vulnerability prioritization. Research has demonstrated that CVSS scores correlate poorly with actual exploitation likelihood [3]. The distribution of CVSS scores is heavily skewed toward higher values, with a substantial percentage of vulnerabilities receiving scores of 7.0 or higher, creating an overwhelming volume of "high priority" issues. Furthermore, CVSS measures theoretical maximum impact rather than real-world exploitation risk, leading to misaligned prioritization efforts.
2.2 Known Exploited Vulnerabilities (KEV) Catalog
The CISA Known Exploited Vulnerabilities catalog represents a paradigm shift toward evidence-based vulnerability management [4]. Launched in 2021 as part of Binding Operational Directive 22-01, KEV documents vulnerabilities with confirmed evidence of active exploitation in the wild. Federal civilian executive branch agencies must remediate KEV-listed vulnerabilities within specified timeframes, typically 2-3 weeks for newly added entries.
KEV inclusion criteria require clear evidence of active exploitation, not merely proof-of-concept code or theoretical exploitability. Vulnerabilities must have clear remediation guidance, typically in the form of vendor-provided patches or specific mitigation instructions. Each entry includes structured data fields: CVE identifier, vendor/project, product name, vulnerability description, required action, and due date for federal compliance.
The catalog’s strength lies in its focus on confirmed threats rather than theoretical risks. Organizations using KEV can confidently prioritize resources toward vulnerabilities that attackers actively exploit. However, KEV has inherent limitations: it is reactive by nature, documenting exploitation only after it occurs; coverage is limited to publicly confirmed exploitation, potentially missing targeted attacks; and the catalog may lag behind rapidly evolving threat landscapes.
@@ 2.3 Exploit Prediction Scoring System (EPSS)
The Exploit Prediction Scoring System represents a data-driven approach to vulnerability prioritization using machine learning [5]. Developed through collaboration between FIRST and academic researchers, EPSS predicts the probability that a vulnerability will be exploited in the wild within the next 30 days. EPSS scores range from 0 to 1 (0-100%), with higher scores indicating greater exploitation likelihood.
The EPSS model incorporates diverse features including CVE-specific characteristics extracted from vulnerability descriptions and technical details, temporal features such as time since publication and day-of-week effects, and threat intelligence signals from multiple commercial and open-source feeds. The model is retrained regularly to maintain prediction accuracy as the threat landscape evolves. Published research demonstrates that EPSS significantly outperforms CVSS for vulnerability prioritization, achieving superior efficiency (proportion of prioritized vulnerabilities that are exploited) and coverage (proportion of exploited vulnerabilities that are prioritized) [6].
EPSS provides daily score updates reflecting the dynamic nature of exploitation risk. Organizations can access current scores through APIs or bulk data downloads, enabling integration with existing vulnerability management workflows. However, as a predictive model, EPSS has inherent uncertainty with both false positives and false negatives. Scores may also exhibit volatility as new threat intelligence emerges or model updates occur.
2.4 Related Research
Several research streams inform our integrated vulnerability management framework, spanning empirical studies on exploitation patterns, alternative prioritization approaches, machine learning techniques, and decision analysis frameworks.
2.4.1 Empirical Studies on Vulnerability Exploitation
Allodi and Massacci [7] provided foundational evidence that only a small fraction of published vulnerabilities are exploited in practice, establishing the theoretical basis for risk-based prioritization. This critical finding has been consistently validated and motivates focusing on vulnerabilities likely to be exploited rather than attempting comprehensive remediation. Building on this insight, Jacobs et al. [8] developed models for improving vulnerability remediation through better exploit prediction, demonstrating that data-driven approaches can significantly outperform traditional severity-based prioritization. Other empirical studies such as Bozorgi et al. [9] and Frei et al. [10] have further explored vulnerability lifecycle patterns and exploitation characteristics, reinforcing the need for risk-based approaches.
2.4.2 Alternative Prioritization Frameworks
The limitations of CVSS have motivated several alternative frameworks. Spring et al. analyzed CVSS limitations extensively [3], with Howland [11] declaring CVSS "ubiquitous and broken." The Stakeholder-Specific Vulnerability Categorization (SSVC) framework, initially proposed in 2019 [12] and refined in 2021 [13], provides context-dependent prioritization based on stakeholder roles and decision contexts. While SSVC offers valuable insights about the importance of organizational context, it requires manual assessment of each vulnerability against multiple decision points, limiting scalability for organizations managing thousands of vulnerabilities.
The Exploit Prediction Scoring System (EPSS), first introduced in 2019 [14] and subsequently refined [5, 6], has emerged as the most successful data-driven alternative to CVSS. Using machine learning to predict exploitation probability, EPSS achieves superior efficiency and coverage compared to traditional approaches. Commercial vendors have also developed proprietary frameworks, including Tenable’s Vulnerability Priority Rating [15] and systems like Vulcon [16], though these often lack transparency or require expensive licenses. The recent releases of EPSS v4.0 and NIST’s Likely Exploited Vulnerabilities (LEV) metric [17], alongside CVSS v4.0 [18], further validate the evolution toward sophisticated predictive methodologies. [^1] This study utilizes EPSS v3.0. The emergence of EPSS v4.0 and LEV [17] reinforces the value of our integration framework, which can readily incorporate such improvements without architectural changes.
2.4.3 Machine Learning and AI Approaches
Researchers have explored various machine learning approaches for vulnerability assessment and prediction. Sabottke et al. [19] demonstrated that Twitter discussions can predict real-world exploits before they appear in traditional threat intelligence feeds. Yamamoto et al. [20] developed text mining approaches for estimating vulnerability scores based on vulnerability descriptions. Almukaynizi et al. [21] created DARKMENTION, a system that monitors dark web forums to predict enterprise-targeted cyberattacks, with subsequent work by Samtani et al. [22] extending dark web analysis through diachronic graph embedding frameworks.
Integration of diverse threat intelligence sources has been explored by Nunes et al. [23], who developed methods for mining darknet and deepnet sources for proactive threat intelligence. Schaberreiter et al. [24] addressed the critical issue of evaluating trust in cyber threat intelligence sources, highlighting the challenges of integrating heterogeneous data. While these approaches show promise, they often require specialized data collection infrastructure, natural language processing capabilities, or access to non-public data sources that may not be accessible to all organizations.
2.4.4 Multi-Criteria Decision Analysis for Risk Assessment
Multi-criteria decision analysis (MCDA) provides structured methodologies for complex risk assessment problems. In the cybersecurity domain, Fridgen et al. [25] proposed a comprehensive framework incorporating technical, economic, and organizational factors. Wang et al. [26] utilized attack graphs combined with MCDA for measuring security risk of networks.
More recently, Ganin et al. [27] developed a decision-analysis-based approach specifically for cybersecurity that quantifies threat, vulnerability, and consequences through structured criteria. While these MCDA approaches offer comprehensive risk evaluation, they typically require extensive manual configuration, subjective weight assignments, and domain expertise that may limit practical deployment.
2.4.5 Operational Perspectives and Integration Gaps
From an operational perspective, Bulut et al. [28] explored vulnerability prioritization through an offensive security lens, emphasizing the importance of attacker-centric thinking. Industry reports from vendors such as Rapid7 [29] provide practical insights into real-world vulnerability trends. Compliance frameworks like PCI DSS [30] mandate specific vulnerability management practices, while government directives such as CISA’s binding operational directives [31, 32] establish regulatory requirements for vulnerability remediation timelines.
NIST’s risk assessment guidelines [33] provide comprehensive frameworks but focus on general risk management rather than specific vulnerability prioritization methodologies. Despite extensive research on individual vulnerability assessment approaches, limited work exists on systematic integration of multiple intelligence sources. Most research focuses on improving individual components rather than their synergistic combination. Commercial vulnerability management platforms have developed proprietary integration approaches, but these typically lack transparency about their methodologies and require expensive subscriptions, limiting accessibility.
2.4.6 Summary of Research Gaps
Our literature review identifies several critical gaps in existing vulnerability management research:
First, there is a lack of systematic integration frameworks that combine multiple vulnerability intelligence sources. While CVSS, EPSS, and KEV each provide valuable but incomplete perspectives, no existing framework systematically leverages their complementary strengths. Second, limited empirical validation exists for proposed approaches. Most research presents theoretical frameworks or limited case studies rather than comprehensive evaluation using real-world exploitation data across extended time periods. Third, many solutions face accessibility barriers, requiring proprietary data sources, specialized infrastructure, or expertise that may not be available to resource-constrained organizations. Finally, sophisticated approaches often exhibit operational complexity that fails to consider practical implementation within existing vulnerability management workflows.
Our Vulnerability Management Chaining framework addresses these gaps by providing an empirically validated, accessible, and operationally simple methodology for integrating established vulnerability intelligence sources. By using exclusively open-source data and transparent decision logic, we enable organizations of all sizes to implement effective risk-based vulnerability management while maintaining flexibility to incorporate emerging intelligence sources as they mature.
3 Vulnerability Management Chaining Framework
@@
This section presents our integrated vulnerability management framework that systematically combines CVSS, EPSS, and KEV to achieve more effective prioritization than any individual approach.
@@ 3.1 Framework Design Principles
Our framework design is guided by several key principles derived from analysis of current vulnerability management challenges and the complementary nature of existing approaches.
Multi-source intelligence integration recognizes that effective vulnerability management requires multiple perspectives on risk. No single metric or data source provides complete visibility into vulnerability threats. By combining technical severity assessment (CVSS), predictive intelligence (EPSS), and confirmed exploitation evidence (KEV), we capture a more comprehensive view of vulnerability risk that addresses both current threats and emerging risks.
Decision tree structure provides clear, actionable guidance for vulnerability prioritization. Unlike abstract scoring systems that produce continuous values requiring interpretation, our decision tree yields specific categorization and recommended actions. This approach aligns with operational security workflows where teams need unambiguous guidance on which vulnerabilities to address and in what order.
Threat-first evaluation prioritizes actual and predicted exploitation over theoretical severity. Traditional CVSS-based approaches evaluate all vulnerabilities regardless of exploitation likelihood, leading to resource waste on vulnerabilities that pose no real threat. By first filtering for exploitation evidence or likelihood, we focus attention on vulnerabilities that represent actual risk to the organization.
Practical implementation using open-source data ensures broad accessibility. All data sources required for our framework—CVSS scores, EPSS predictions, and KEV catalog membership—are freely available and regularly updated. This approach democratizes advanced vulnerability management capabilities without requiring expensive commercial threat intelligence subscriptions.
3.2 Decision Tree Structure
Figure 1 illustrates our complete Vulnerability Management Chaining decision tree. The framework employs a two-stage evaluation process that first assesses threat likelihood, then evaluates vulnerability characteristics to determine final prioritization.
All CVEs (28,377 total) In KEV Catalog? EPSS $\geq$ 0.088? CVSS $\geq$ 7.0? CVSS $\geq$ 7.0? CISA KEV 922 CVEs EPSS Model ML Predictions CVSS Scores Base Metrics CVSS Scores Base Metrics Critical Priority Immediate Remediation Monitor Track for Changes High Priority Scheduled Remediation Defer Standard Patching YesNoYesNoYesNoNoYes Stage 1 Results:
KEV: 922 CVEs
EPSS $\geq$ 0.088: 2,984 Combined: 3,682
Final Prioritization:
Critical: 632
High: 2,419 Monitor: 631 Defer: 24,695
Figure 1: Vulnerability Management Chaining decision tree framework showing the two-stage evaluation process. Stage 1 applies threat-based filtering using KEV membership or EPSS threshold ($\geq$0.088). Stage 2 uses CVSS scores ($\geq$7.0) for severity assessment to enable informed deprioritization. The framework categorizes vulnerabilities into four priority levels: Critical (immediate remediation), High (scheduled remediation), Monitor (track for changes), and Defer (standard patching).
The first stage identifies vulnerabilities with evidence of exploitation threat. We check KEV catalog membership to identify vulnerabilities with confirmed active exploitation. For vulnerabilities not in KEV, we evaluate EPSS scores against a threshold of 0.088 to identify those with elevated exploitation probability. This threshold represents the intersection of optimal efficiency and coverage based on published EPSS research. Vulnerabilities failing both checks are deferred to standard patching cycles, as they show no evidence of exploitation threat.
The second stage applies vulnerability severity assessment to enable informed deprioritization. For vulnerabilities identified as threats in Stage 1, we evaluate CVSS base scores against a threshold of 7.0 (high severity). This assessment distinguishes between high-impact vulnerabilities requiring urgent attention and lower-impact issues that may warrant different handling despite exploitation evidence.
Our framework yields four distinct prioritization categories. Critical Priority encompasses vulnerabilities in the KEV catalog with CVSS $\geq$ 7.0, representing confirmed exploitation of high-impact vulnerabilities. These demand immediate remediation, typically within days. High Priority includes vulnerabilities with EPSS $\geq$ 0.088 and CVSS $\geq$ 7.0, indicating predicted exploitation of high-impact vulnerabilities. These should be addressed through scheduled remediation, typically within 2-4 weeks. Monitor category covers vulnerabilities with exploitation evidence (KEV or high EPSS) but CVSS < 7.0. These require tracking for changes in threat landscape or impact assessment. Defer category includes vulnerabilities with no exploitation evidence (not in KEV and EPSS < 0.088), which can be addressed through standard patching cycles.